OpenStudy (anonymous):

g(x)=x^2+1 (a) find the average rate of change from -1 to 2. (b) find an equation of the secant line containing (-1, g(-1)) and (2, g(2)). g(x)=x^2+1 (a) find the average rate of change from -1 to 2. (b) find an equation of the secant line containing (-1, g(-1)) and (2, g(2)). @Mathematics

OpenStudy (anonymous):

a) average rate of change = (g(2)-g(-1))/(2-(-1)) = (5-2)/(2+1) = 1 b) the equation is (y-g(2)) = 1*(x-2), i.e. y-5 = x-2, i.e. y = x+3

OpenStudy (anonymous):

THANK YOU!!!